# Human-Object Interaction with Vision-Language Model Guided Relative Movement Dynamics

### v1

@@ 5 Conclusion
In this paper, we proposed a unified framework for Human-Object Interaction (HOI) that leverages the concept of Relative Movement Dynamics (RMD) to model complex spatial-temporal interactions between humans and objects. By integrating vision-language models (VLMs) for high-level planning and goal-conditioned reinforcement learning for precise task execution, our approach enables long-term, physically plausible interactions. We also introduced InterPlay, a diverse dataset designed to support training and evaluation across a wide range of static and dynamic HOI tasks. Extensive experiments demonstrate that our framework outperforms state-of-the-art methods in both single-task and multi-task scenarios, achieving higher task completion rates and greater interaction precision.
@@ 6 Acknowledgement
This work was supported by Shanghai Local College Capacity Building Program (23010503100), National Natural Science Foundation of China (No.62406195), MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration, HPC Platform of ShanghaiTech University and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
@@
Appendix A Demo Video
To complement the qualitative results discussed in the main paper, we provide several demonstration videos that offer detailed visualizations of our method. These videos not only serve to illustrate the qualitative comparisons with competing approaches but also reinforce the effectiveness of our framework across various real-world scenarios. For a comprehensive understanding, we encourage the reader to refer to the supplementary materials accompanying this paper.
Single Task Scenario. Demo4.mp4 and Demo5.mp4 present a qualitative comparison between our method and UniHSI [45] for the sitting task. In these videos, when the agent attempts to stand up from a seated position, we observe that the method employed by UniHSI relies excessively on high-frequency jitter to twist the pelvis. This excessive reliance causes the agent to become stuck in a local minimum of the goal-reaching objective, leading to unrealistic and jerky motion that deviates from natural human behavior. In contrast, our method generates smooth, human-like motion, ensuring a natural transition from sitting to standing and more faithfully replicating realistic human behavior.
Additionally, Demo6.mp4 and Demo7.mp4 compare our approach with InterPhys [9] in the door-opening task. Although InterPhys achieves strong quantitative results, its generated motion appears unrealistic. Specifically, during the door-opening task, InterPhys uses only its body to push the door open, resulting in an unnatural movement pattern. In contrast, our method employs a coordinated use of both hands, which produces a more lifelike interaction and aligns closely with typical human behavior observed in similar scenarios.
Multi Task Scenario. Demo1.mp4, Demo2.mp4, and Demo3.mp4 further demonstrate long-horizon interactions involving static, dynamic, and articulated objects within a realistic indoor furniture setting. These videos illustrate how our method robustly handles multi-task scenarios by leveraging three key strengths of our framework:
  1. 1.
    Human-like Motion: Our approach is designed to ensure that the agent avoids undesirable states while transitioning smoothly between tasks. This results in fluid and coherent motions that are consistent with natural human biomechanics.
  2. 2.
    Unified Objective Design: By employing a consistent objective design across various interaction types, our framework effectively simplifies the task of multi-task learning. This unified approach reduces the need for extensive tuning of separate objectives for different tasks.
  3. 3.
    Vision-Language Model Capabilities: The integrated Vision-Language Model (VLM) leverages its visual understanding of the spatial layout along with rich world knowledge to generate plausible and contextually relevant action plans. This capability is crucial for adapting to complex environments and ensuring realistic task execution.
Figure 4: An overview of RMD Planner pipeline.
Appendix B Details of RMD Planner
As shown in Fig. 4, receiving the top-view image of the surrounding scene context and the text instruction, the RMD Planner outputs sequential sub-step plans in the form of structured JSON form so that the output can be processed directly using python script. To ensure the RMD Planner can work as we expect, we utilize different sections of prompt while each section supports individual functions. The detailed prompt for each section is listed below:
Figure 5: Details of prompt section 1.
Scene-context Understanding. Figure 5 provides the detailed prompt which guides the planner. Note that we simplify the spatial relationships between objects into seven types: center, forward, back, left, right, up, and down, such relative orientation is defined according to the sub-coordinate system of the reference object.
Figure 6: Details of prompt section 2.
RMD Concept Definition. In Fig. 6, we introduce the concept of Relative Movement Dynamics, represented as a bipartite graph modeling interactions between human body parts and object parts over time. The graph includes nodes for human and object parts, with edges capturing relative movement trends, such as stationary, approaching, receding, or unstable dynamics.
Figure 7: Details of prompt section 3.
Specific Plan Instance. Figure 7 illustrates an instance of the output plan, demonstrating the Relative Movement Dynamics during a scenario where a humanoid bends down to grab a box with both hands. The graph details specific relationships and movement trends between body parts and object parts, showcasing the concept in action. Besides, it also plans the target position of the humanoid root and the target position of the object root.
Figure 8: Details of prompt section 4.
@@ Appendix B Details of RMD Planner
Idea Outline. Fig. 8 provides an overview of our approach, highlighting the expected input and output of the interaction planner. Human-object interaction tasks are broken into sub-sequences, each characterized by consistent movement dynamics until transitioning to the next stage.
Figure 9: Details of prompt section 5.
@@
Plan Rules. Figure 9 outlines the rules that the planner must follow, underscoring their importance in prompt engineering. These rules ensure that the planner generates accurate and coherent sub-sequences by providing clear guidelines on task interactions, sequence formatting, and criteria for determining movement dynamics. By adhering to these rules, the planner effectively breaks down complex interactions into manageable steps, enhancing the quality and reliability of the generated plans.
Figure 10: Details of prompt section 6.
Figure 11: Details of prompt section 6.
Reference Example. Figures 10, 11 and 12 sequentially presents a complete example of both input and output, clearly illustrating the concept and workflow of the planner.
@@ Appendix C Additional Details of Reinforcement Learning
Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. Further details on the definitions of state, reward, as well as the training procedure, are provided below.
State Definition. The state is divided into two components: the proprioception of the humanoid and object, and the goal condition. The proprioception component forms a 276-dimensional vector, encompassing the position, rotation, velocity, and angular velocity of the humanoid joints, as well as the bounding box of the object. While the root height is measured in the global reference frame, all other components are defined relative to the local frame of the character. The goal condition includes a 9×9 ego-centric heightmap, the target positions of the humanoid and object, kinematic relation encoding, and dynamic encoding for each part, resulting in a 222-dimensional vector.
Reward Definition. Recall the RMD reward $r_{\text{RMD}}$ at time step $t$ is the weighted sum of all the parts.
$r_{\text{RMD}}=\sum_{k=1}^{15}w_{k}\cdot r_{RMD}^{k}.$ (12)
The RMD reward for the $k$-th part $r_{RMD}^{k}$ is defined as:
$r_{RMD}^{k}=\begin{cases}r_{rv}^{\text{still}},&\text{if }v_{ij}=0\\ w_{pos}\cdot r_{rp}^{\text{close}}+w_{vel}\cdot r_{rv}^{\text{close}},&\text{if }v_{ij}=1\\ w_{pos}\cdot r_{rp}^{\text{away}}+w_{vel}\cdot r_{rv}^{\text{away}},&\text{if }v_{ij}=2\\ 1,&\text{if }v_{ij}=3\end{cases}$ (13)
The term $r_{rv}^{\text{still}}$ is defined by minimizing the dot product between the relative position and relative velocity of two particles. This term encourages either the relative velocity between the two particles to approach zero or the angle between their relative velocity and relative position to approach $\frac{\pi}{2}$.
$r_{rv}^{\text{still}}=\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\cdot\tilde{\mathbf{v}}^{ij}_{t}\right\rVert_{2}\right)$ (14)
The term $r_{rp}^{\text{close}}$ is defined to minimize the $\ell_{2}$ norm of the relative position between the two particles, thereby encouraging them to move closer.
@@
$r_{rp}^{\text{close}}=\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert_{2}\right)$ (15)
The term $r_{rv}^{\text{close}}$ encourages the projection of the relative velocity onto the direction of the relative position to approach a threshold. Note that the directions of these vectors are always opposite.
@@
$r_{rv}^{\text{close}}=\exp\left(-\left\lVert\tilde{\mathbf{v}}^{ij}_{t}\cdot\frac{\tilde{\mathbf{p}}^{ij}_{t}}{\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert}+\epsilon\right\rVert_{2}\right)$ (16)
The term $r_{rp}^{\text{away}}$ is defined to maximize the $\ell_{2}$ norm of the relative position between the two particles, promoting a greater separation between them.
$r_{rp}^{\text{away}}=1-\exp\left(-\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert_{2}\right)$ (17)
The term $r_{rv}^{\text{away}}$ encourages the projection of the relative velocity onto the direction of the relative position to approach a threshold. Here, the directions of these vectors are always aligned.
$r_{rv}^{\text{away}}=\exp\left(-\left\lVert\tilde{\mathbf{v}}^{ij}_{t}\cdot\frac{\tilde{\mathbf{p}}^{ij}_{t}}{\left\lVert\tilde{\mathbf{p}}^{ij}_{t}\right\rVert}-\epsilon\right\rVert_{2}\right)$ (18)
Figure 12: Details of prompt section 6.
Figure 13: Visualization of long-term interaction with objects in an indoor home setting (part 1).
Figure 14: Visualization of long-term interaction with objects in an indoor home setting (part 2).
Figure 15: Visualization of long-term interaction with objects in an indoor home setting (part 3).
Figure 16: Visualization of long-term interaction with objects in an indoor home setting (part 4).
Figure 17: Visualization of long-term interaction with objects in an indoor home setting (part 5).
Figure 18: Visualization of long-term interaction with objects in an indoor home setting (part 6).
Figure 19: Visualization of long-term interaction with objects in an indoor home setting (part 7).
Figure 20: Visualization of long-term interaction with objects in an indoor home setting (part 8).

### v2

@@ 5 Conclusion
In this paper, we present a physics-based framework for synthesizing diverse Human-Object Interactions, guided by VLMs, through our proposed Relative Movement Dynamics —a spatiotemporal representation that structures part-level relationships between humans and objects. RMD automates the generation of goal states and reward functions for reinforcement learning, eliminating the need for manual engineering while supporting static, dynamic, and articulated interactions. Our method bridges high-level semantic reasoning with low-level physics-based control, producing natural motions that outperform task-specific and LLM-based baselines in both realism and adaptability. To facilitate evaluation, we introduce the Interplay dataset for long-horizon HOI tasks. Extensive results demonstrate the effectiveness of our approach on both single-task and long-horizon multi-task scenarios. Future directions include integrating diffusion models to improve motion diversity and exploring joint training of VLMs with physics-based controllers to enhance generalization and control fidelity.
@@ 6 Acknowledgement
This work was supported by National Natural Science Foundation of China (62406195, 62303319), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, Core Facility Platform of Computer Science and Communication of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), and Shanghai Engineering Research Center of Intelligent Vision and Imaging.
@@
[…]
Appendix
Figure A1: An overview of VLM-guided RMD Planner pipeline.
Appendix A Details of RMD Planner
As shown in Fig. A1, given a top-view image of the surrounding scene and a textual instruction, the VLM-guided RMD Planner generates a sequence of sub-step plans in the form of structured JSON. This output format facilitates direct downstream processing via Python scripts.
To ensure that the planner performs as intended, we design prompts that explicitly guide it to understand the concept of Relative Movement Dynamics (RMD), perceive the contextual scene information, decompose high-level instructions, and produce outputs in a structured format consistent with our framework. For clarity and modularity, we organize the prompt into distinct sections, each fulfilling a specific functional role as illustrated in Fig. A2 and Fig. A3. The content of each section is detailed below.
Scene-context Understanding. This section of the prompt helps the planner interpret the scene context and determine the target destination using a simplified relative spatial relationship. Specifically, we reduce spatial relations to seven canonical types: center, forward, back, left, right, up, and down.
Figure A2: Details of prompt section.
Figure A3: Details of prompt section.
These relative positions are defined with respect to a local coordinate system centered on the reference object.
RMD Concept Defination. This section of the prompt helps the planner understand the concept of RMD, clarifies its intended semantics, introduces a method for instantiating it, and provides a formal definition of its instantiated form.
Specific Plan Instance. This section presents a templated instance of an RMD-based plan, illustrating the correspondence between the template elements and RMD components, as well as how such an instance aligns with a real-world interaction scenario. This helps the planner internalize the concept and encourages outputs that conform to the desired format.
@@ Appendix A Details of RMD Planner
Idea Outline. This section offers a high-level overview of our approach, specifying the expected input and output of the interaction planner. Human–object interaction tasks are decomposed into a sequence of sub-steps, each characterized by consistent movement dynamics.
@@
Plan Rules. This section enumerates the rules that the planner must follow, emphasizing their significance in prompt engineering. These rules define task interaction boundaries, output formatting, and criteria for determining movement dynamics. By adhering to these rules, the planner can effectively decompose complex interactions into manageable and coherent sub-sequences.
Reference Example. Empirical findings from the NLP community indicate that one-shot prompting often yields significantly better results than zero-shot prompting. Guided by this insight, we include a complete input–output example in the prompt to help the planner better understand the task expectations and output format.
Figure A4: Visualization of long-term interaction with objects in an indoor home setting (part 1).
Figure A5: Visualization of long-term interaction with objects in an indoor home setting (part 2).
Figure A6: Visualization of long-term interaction with objects in an indoor home setting (part 3).
Figure A7: Visualization of long-term interaction with objects in an indoor home setting (part 4).
Figure A8: Visualization of long-term interaction with objects in an indoor home setting (part 5).
Appendix B InterPlay Dataset
To evaluate our framework in long-horizon, multi-task scenarios, we construct a novel dataset, InterPlay, which encompasses both static and dynamic interaction tasks under diverse indoor scene contexts. Each data sample includes: (i) a list of processed 3D assets ready for direct use in simulation, (ii) the corresponding 3D scene layout, (iii) a top-view rendered image, and (iv) a natural language instruction describing the intended task.
In building the InterPlay dataset, we incorporate high-quality 3D object assets from several existing datasets, including PartNet [62], 3D-FRONT [63], SAMP [27], CIRCLE [64], and PartNet-Mobility [65]. From these sources, we carefully select 117 assets spanning categories such as bed, chair, sofa, washing machine, box, door, window, and wardrobe. Each asset is normalized, segmented into functional parts, and associated with part-level point clouds, based either on existing annotations or manual labeling.
To generate scene layouts, we apply a set of heuristic rules that ensure: (i) a well-distributed arrangement without object collisions, (ii) semantically meaningful placements aligned with typical household settings, and (iii) a minimum of two interactive objects and at least two furniture items per scene, thereby enriching the contextual complexity of human–object interaction. We then populate these layout templates by randomly sampling from the processed asset pool.
For each completed layout, we render a top-view image and employ GPT-4V [60] to generate a corresponding task instruction grounded in the scene context. Finally, we apply our VLM-guided RMD Planner to produce interaction plans, which are manually reviewed and refined before inclusion in the dataset.
In total, InterPlay contains 1,210 interaction plans, covering a wide range of HOI behaviors involving static, dynamic, and articulated objects with varying temporal lengths, all situated in realistic, household-inspired environments.
@@ Appendix C Additional Details of Reinforcement Learning
Our physics-based animation framework is built upon goal-conditioned reinforcement learning. At each time step $t$, the agent samples an action from its policy $\pi(a_{t}\mid\mathbf{s}_{t},\mathbf{g}_{t})$ based on the current state $\mathbf{s}_{t}$ and the goal state $\mathbf{g}_{t}$. After executing the action, the environment transitions to the next state $\mathbf{s}_{t+1}$, and the agent receives a task reward $r^{G}(\mathbf{s}_{t},\mathbf{g}_{t},\mathbf{s}_{t+1})$. Further details are provided below.
Action. We use a simulated humanoid character with 28 degrees of freedom (DoF), represented as a kinematic structure composed of 15 rigid bodies. The action space is defined as the target joint positions for 28 proportional–derivative (PD) controllers. At each timestep, the predicted actions specify these targets, which are then translated by the PD controllers into joint torques to drive the character’s motion.
Proprioception. The proprioceptive state $\mathbf{s}_{t}$ is a 223-dimensional feature vector that encodes the character’s physical state. It includes the positions, rotations, linear velocities, and angular velocities of all rigid bodies, expressed in the character’s local coordinate frame. The only exception is the root height, which is represented in the world coordinate frame to preserve global elevation information.
Task Reward. Recall that the term $r_{\text{RMD}}$ encourages each edge $e_{ij}\in\mathcal{B}$, representing a (human-part, object-part) pair, to follow the relative motion dynamics specified by its associated edge weight $w_{ij}$, as planned by the VLM. Specifically, $r_{\text{RMD}}$ is computed by summing the individual alignment reward $r_{\text{rmd}}^{ij}(\cdot)$, each of which measures how well the pair $(p_{h_{i}},p_{o_{j}})$ follows the motion pattern defined by $w_{ij}$, weighted by a coefficient $\lambda_{ij}$,
$r_{\text{RMD}}=\sum_{(i,j)\in E}\lambda_{ij}\cdot r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right).$ (15)
@@
Specifically, we formally define $r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right)$ as,
@@
$r_{\text{rmd}}\!\left(\tilde{\mathbf{p}}_{ij},\tilde{\mathbf{v}}_{ij},w_{ij}\right)=\begin{dcases}\exp\left(-\left(\tilde{\mathbf{p}}_{ij}\cdot\tilde{\mathbf{v}}_{ij}\right)^{2}\right),&w_{ij}=0,\\[4.0pt] \frac{1}{2}\exp\left(-\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}^{2}\right)+\frac{1}{2}\exp\left(-\left(\tilde{\mathbf{v}}_{ij}\cdot\frac{\tilde{\mathbf{p}}_{ij}}{\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}}+v^{\star}\right)^{2}\right),&w_{ij}=1,\\[4.0pt] \frac{1}{2}\left(1-\exp\left(-\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}^{2}\right)\right)+\frac{1}{2}\exp\left(-\left(\tilde{\mathbf{v}}_{ij}\cdot\frac{\tilde{\mathbf{p}}_{ij}}{\left\lVert\tilde{\mathbf{p}}_{ij}\right\rVert_{2}}-v^{\star}\right)^{2}\right),&w_{ij}=2,\\[4.0pt] 1,&w_{ij}=3.\end{dcases}$
The behavior of each reward term varies depending on the value of $w_{ij}$:
- When $w_{ij}=0$, the reward promotes either a vanishing relative velocity or an orthogonal orientation between the relative position and velocity, thereby minimizing directional alignment.
- For $w_{ij}=1$, the goal is to maintain spatial proximity while aligning the relative velocity with the direction of the relative position. In particular, the velocity projection is encouraged to approach a predefined target $v^{\star}=1\,\mathrm{m/s}$.
- When $w_{ij}=2$, the reward favors increasing separation between parts, while ensuring that the leaving speed along the direction of displacement remains close to $v^{\star}$.
- Finally, $w_{ij}=3$ denotes an unstable or unpredictable relative motion trend. Here, the reward is set to a neutral constant value of 1 to avoid imposing misleading gradients.
Training details in a single-task scenario. Our policy network is divided into a critic network and an actor network, each starting with three multilayer perceptron (MLP) layers configured with [1024, 1024, 512] units. The discriminator network follows a similar architecture, also consisting of three MLP layers with [1024, 1024, 512] units. The hyperparameters used during the training process are detailed Tab. A1,
Table A1: Hyperparameters for RMD.
Hyperparameter Value Hyperparameter Value Hyperparameter Value
Num. envs 4096 Max episode length 450 Num. epochs 25,000
Discount factor $\gamma$ 0.99 GAE parameter $\lambda$ 0.95 Minibatch size 16384
Optimizer Adam Learning rate 2e-5 Gradient clip norm 1.0
PPO horizon length 32 PPO clip $\epsilon$ 0.2 PPO miniepochs 6
Actor loss weight 1 Critic loss weight 5 Discriminator loss weight 5
AMP batch size 512 AMP sequence length $t^{*}$ 10 AMP grad. penalty $w^{gp}$ 5
Task reward weight $w^{G}$ 0.5 Disc. reward weight $w^{S}$ 0.5 Bounds loss coefficient 10
Some baseline methods either did not implement certain HOI tasks (e.g., open, push) or failed to account for post-interaction transitions with static objects (e.g., sit, lie). To enable a fair and thorough comparison, we re-implemented these baselines with carefully hand-crafted goal states and reward functions for each task. This ensures that we can comprehensively evaluate the full interaction cycle: approaching the object, interacting with it, and departing afterward, across a range of individual HOI tasks. All methods were trained using a shared set of hyperparameters to maintain consistency.
Training details in a long-horizon multi-task scenario. Our policy consists of separate actor and critic networks, each implemented as a three-layer multilayer perceptron (MLP) with hidden dimensions of [1024, 1024, 512]. The discriminator network adopts the same architecture. Both our method and the baselines are trained using a two-stage procedure:
(i) Single-task pre-training. We divide the parallel simulation environments into distinct subsets according to task type and difficulty, allowing the agent to learn basic interaction skills in a focused manner.
(ii) Long-horizon multi-task post-training. For each environment instance, tasks are randomly sampled from our InterPlay dataset, promoting the agent’s ability to handle long-horizon human–object interaction tasks in complex and diverse scenes.
To extend InterPhys to a multi-task setup, we concatenate all possible goal states and use a one-hot encoding to indicate the task identity. Additionally, we implement a conversion script to translate the sequential plans in our dataset into the format required by other baselines (i.e., sequential goal states), thereby unifying the training pipeline. During execution, a rule-based finite-state machine is used to manage task transitions in both InterPhys and TokenHSI.
Appendix D Visualizations in Supplementary Material
To complement the qualitative results presented in the main paper, we provide a demonstration video that combines the key aspects of our method. This video offers detailed visualizations showcasing the effectiveness of our framework in various real-world scenarios and compares it with competing approaches. It includes an overview of the motivation, challenges, main pipeline, and both quantitative and qualitative results. Additionally, the video highlights long-horizon interactions in both single-task and multi-task scenarios, involving static, dynamic, and articulated objects within a realistic indoor setting. In the single-task scenario, we present a comparison of our method with UniHSI [13] for the sitting task and InterPhys [15] for the door-opening task. While UniHSI struggles with high-frequency jitter and unrealistic motion, our method ensures smooth, human-like transitions. Similarly, in the door-opening task, our approach uses coordinated hand movements for a more lifelike interaction, in contrast to InterPhys, which relies solely on body movement and produces unnatural results. The multi-task scenario illustrates how our method excels by leveraging human-like motion, a unified objective design across different tasks, and the capabilities of Vision-Language Models to generate contextually relevant action plans. Additionnaly, we provide detailed visualizations alongside the Planner’s output, as shown in Fig. A4, Fig. A5, Fig. A6, Fig. A7 and Fig. A8.
Appendix E Limitations and Future Work
Limitations. While our framework advances dynamic human-object interaction modeling, it currently focuses on single-agent scenarios and does not address multi-agent interactions (e.g. collaborative tasks involving multiple humans or agents) or social dynamics (e.g. social impact considerations in shared environments). This restricts its applicability to real-world settings requiring coordination among multiple agents, such as collaborative kitchens or assistive care scenarios. Additionally, the current VLM-based planner may struggle with long-horizon tasks that demand extended temporal reasoning and hierarchical planning. For instance, tasks like cooking multi-step meals or tidying interconnected household spaces remain challenging due to the lack of explicit task decomposition mechanisms.
Future Work. To address these gaps, future research directions include extending the framework to multi-agent reinforcement learning, where agents could learn adaptive communication protocols and dynamic role allocation in socially rich environments. Simultaneously, integrating chain-of-thought prompting into the VLM-based planner could enhance its capacity for stepwise task decomposition (e.g., sequencing ingredient preparation, cooking, and serving), improving robustness in household robotics scenarios. Combining multi-agent social coordination and hierarchical reasoning would bridge the gap between atomic interactions and goal-oriented task execution, enabling scalable deployment in collaborative human environments.
Appendix F Broader Social Impact
Our method can be used to generate realistic human–object interactions for virtual avatars, simulate ergonomic behaviors, or control assistive humanoid robots in household environments. By reducing the need for hand-crafted motion design, it can accelerate the development of home robotics, virtual animation, and simulation tools. Additionally, it has the potential to improve accessibility in AR/VR applications by enabling more natural and intuitive interaction modeling, contributing to positive social impact in healthcare, rehabilitation, and inclusive technology.